ganeshie8 (ganeshie8):

Prove/Disprove \(\large 2^{p − 1} \equiv 1 \mod p^2\) has no solutions \(p\) is a prime number

OpenStudy (dangerousjesse):

:)

ganeshie8 (ganeshie8):

I am not able to find solutions for your first equation itself haha! looks the proof would be interesting xD

OpenStudy (dangerousjesse):

I agree, it's a fairly mind-boggling equation :P

OpenStudy (ikram002p):

hmm 2^p-1=1 mod p^2 lets assume it has a solution , then 2^(p-1/2)= 1 mod p hmm lets see any contradiction

OpenStudy (dangerousjesse):

So, basically, \(\large p=\frac{1}{2}\)

OpenStudy (ikram002p):

no there is not such a thing , im looking for Elulers criterion it should contradict this thing , but not sure if we can apply 2

OpenStudy (ikram002p):

only found primitive root :(

ganeshie8 (ganeshie8):

if it helps, there are no solutions for the first 1000 primes : http://primes.utm.edu/lists/small/1000.txt

ganeshie8 (ganeshie8):

however the equation \(\large \color{Red}{3}^{p − 1} \equiv 1 \mod p^2\) has a solution : \(\large p = 11\)

ganeshie8 (ganeshie8):

brb...

OpenStudy (ikram002p):

not in general , it says :- if p is an odd prime then there exist a primitive root of p , such that \(r^p-1 \neq 1 \mod p^2\) hmm idk for 2 :(

OpenStudy (ikram002p):

ok lol

OpenStudy (ikram002p):

i just read it for k>=3 the integer 2^k has no primitive roots :3 done

OpenStudy (ikram002p):

@ganeshie8 how did you check 1000 numbers ?

ganeshie8 (ganeshie8):

http://ideone.com/a1fgEh scroll down

ganeshie8 (ganeshie8):

that shows remainders : \(\large (2^{p-1}-1) \mod p^2\)

OpenStudy (fibonaccichick666):

could use that number theory thingy, let me grab my book

ganeshie8 (ganeshie8):

sure :) i have exhausted all the trials... not getting any new ideas >.<

OpenStudy (fibonaccichick666):

ok so could we show that \(2^{p-1}\) is not relatively prime since 1 isn't a prime? and so 1 will always go into p^2?

OpenStudy (fibonaccichick666):

so it fails the requirement?

ganeshie8 (ganeshie8):

you want to factor is it ? looks like a good idea !

ganeshie8 (ganeshie8):

\(\large 2^{p-1} - 1^{p-1} = (2-1)(2^{p-2 } + \cdots + 1)\)

ganeshie8 (ganeshie8):

hmm~

OpenStudy (fibonaccichick666):

so there is a thm that states g is a prim root mod m then g^r is a prim rootmod m iff \(gcd(r, \phi{m})=1\)

OpenStudy (fibonaccichick666):

could we extrapolate from there?

OpenStudy (fibonaccichick666):

or since 2^p-1=1modp maybe we could use that?

ganeshie8 (ganeshie8):

leme pull up my NT notes :) don't seem to remember a thing lol

OpenStudy (fibonaccichick666):

that's what I'm using... i had a polish teacher.

OpenStudy (fibonaccichick666):

that class was awful

OpenStudy (fibonaccichick666):

AH here so if we can prove that the only powers of 2 with a prim root are 2 and 4 we should be able to say the only possible ones for p squared are 4 and 16 so the only possible primes are 3 and 17

OpenStudy (fibonaccichick666):

5 and 17 but i don't think that's right

OpenStudy (fibonaccichick666):

5 would be 2^4=16/=1mod25 17 would be 2^16=65536/=1MOD289

OpenStudy (fibonaccichick666):

nope ok so, we can prove that the only integers that can have prim roots are 2,4,p^r, and 2p^r st p is odd

OpenStudy (fibonaccichick666):

Maybe this?

OpenStudy (fibonaccichick666):

This first

OpenStudy (fibonaccichick666):

maybe one of the exercises can lead us to something

OpenStudy (ikram002p):

isit burton book /.?

OpenStudy (fibonaccichick666):

number 5 is espec. interesting

OpenStudy (fibonaccichick666):

andrews

OpenStudy (fibonaccichick666):

it's from dover publishing

OpenStudy (ikram002p):

its also chapter 7 to me >.< same questions xD

OpenStudy (fibonaccichick666):

lol

OpenStudy (fibonaccichick666):

it's evilll isn't it?

OpenStudy (ikram002p):

yeah ,hehe

OpenStudy (fibonaccichick666):

but #5 i think could be extremely useful here

OpenStudy (ikram002p):

brb

OpenStudy (fibonaccichick666):

it intentionally leaves out p^2

OpenStudy (fibonaccichick666):

but if we swing over to abstract algebra and thing about cyclic groups, we are essintially saying that this has no generators

OpenStudy (fibonaccichick666):

hmmm